# SnapKV [Uncaptioned image] : LLM Knows What You are Looking for Before Generation

### v1

@@ Abstract
We discover that each attention head in the model consistently focuses on specific prompt attention features during generation. Meanwhile, this robust pattern can be obtained from an ‘observation’ window located at the end of the prompts. Drawing on this insight, SnapKV automatically compresses KV caches by selecting clustered important KV positions for each attention head. Our approach significantly reduces the growing computational overhead and memory footprint when processing long input sequences. Specifically, SnapKV achieves a consistent decoding speed with a 3.6x increase in generation speed and an 8.2x enhancement in memory efficiency compared to baseline when processing inputs of 16K tokens. At the same time, it maintains comparable performance to baseline models across 16 long sequence datasets. Moreover, SnapKV can process up to 380K context tokens on a single A100-80GB GPU using HuggingFace implementation with minor changes, exhibiting only a negligible accuracy drop in the Needle-in-a-Haystack test. Further comprehensive studies suggest SnapKV’s potential for practical applications. Our code is available at  https://github.com/FasterDecoding/SnapKV.
@@ 1 Introduction
Figure 1: The graph shows the simplified workflow of SnapKV, where the orange area represents the group of positions per head clustered and selected by SnapKV. These clustered features are then used to form a new Key-Value pair concatenated with the tokens in the observation window (denoted as ‘Window’). Together, the selected prefix and observation windows constitute the new KV cache utilized for the generation.
Many inspiring works have successfully expanded LLMs to handle longer contexts, overcoming the difficulties in context maintenance and attention mechanism scalability, such as GPT-4 [1] and Command-R [2] with context length 128K, Claude-3 [3] with 200K, and Gemini-Pro-1.5 with 1M [4]. Despite their impressive capabilities, LLMs still face significant challenges when dealing with long context inputs. Specifically, the KV caches in attention calculation become an obstacle in efficiently processing long context. During inference time, as input length increases, the decoding speed per step grows linearly due to the computation for attention across past KVs. Moreover, the large KV cache created during prompting requires significant on-chip and off-chip memory, increasing hardware demands and limiting model scalability.
There are many perspectives to mitigate these problems, including KV cache eviction during token generation [5, 6, 7, 8]. However, most of these methods lack a detailed evaluation of the generated context in a long-context setting. Moreover, they mainly focus on optimizing the KV cache appended during generation steps, while overlooking the realistic problem of compressing KV cache for input sequences, which is typically the bottleneck in memory efficiency. In practical applications such as chatbots and agents, where inputs can be multi-turn conversations, extensive articles or codebases [1, 9, 10], input sizes are often much larger than the sizes of generated responses, resulting in significant overhead. Additional challenge lies in compressing such vast inputs without losing crucial information for accurate generation, especially in scenarios with various noisy contexts.
In our paper, we identify the patterns of these important prompt attention features during generation. To validate the robustness of this finding, we also design a thorough set of experiments across diverse inputs in terms of length, format, and content. Based on our observations, we derive an innovative and intuitive method, SnapKV, which can effectively compress the KV cache for long sequence inputs without compromising the model’s accuracy. Our contributions are as follows:
  • We design experiments to explore the patterns of attention features in output generation, focusing on three key questions:
@@
      Is there a consistent pattern in the attention allocated to prompt tokens?
@@
      How does the context and instruction positioning influence this attention allocation pattern?
    3. 3.
      Does the nature of the user’s instructions play a role in shaping these attention patterns?
    Our finding suggests that most of the LLMs’ attention allocation of input sequence remains unchanged during generation. Thus, LLMs knows what you are looking for before generation.
  • We develop an efficient algorithm, SnapKV, inspired and validated by extensive observations and testing. SnapKV intelligently identifies important KVs with minimal modification (See Fig. 1). The algorithm can be easily integrated into popular deep-learning frameworks with just a few code adjustments.
  • We evaluate SnapKV for accuracy and efficiency across diverse LLMs and long-sequence datasets, affirming its improvement over previous work and comparability to conventional KV caching. Furthermore, we conduct the Needle-in-a-Haystack test to demonstrate its memory efficiency and illustrate decoding speed enhancements through varied batch sizes and input lengths. In addition, SnapKV’s integration with a leading RAG model showcases its extended performance capabilities. We also show that SnapKV can be combined orthogonally with other acceleration strategies such as parallel decoding.
@@ 2 Related Works
Many previous works address the KV cache compression by evicting the KV cache using different algorithms. For example, StreamLLM [5] maintains the first few tokens and the local tokens to effectively reduce the KV cache size. However, it faces the challenge of losing important information since it continuously evicts the KV cache. [^1] https://github.com/mit-han-lab/streaming-llm?tab=readme-ov-file#faq Another perspective is to compress the KV cache for generation steps. Heavy-Hitter Oracle [6] introduces a KV cache eviction policy that greedily selects tokens during generation steps based on a scoring function derived from cumulative attention. While this approach effectively compresses the KV cache for generated tokens, it overlooks compression of the input sequence KV cache, which is crucial for reducing memory and computational overhead. Building on a similar concept, Adaptive KV Compression (FastGen) [8] implements a dual-phase algorithm that encompasses four KV cache compression policies. Initially, it identifies optimal policies through profiling results obtained from prompt encoding. Subsequently, it dynamically evicts caches during the generation phase based on these policies. Nonetheless, it faces the similar problem with H2O. ScissorHands [7] focuses on identifying and retaining pivotal tokens that exhibit a consistent attention weight pattern with previous token windows during generation steps. However, this method concentrates solely on the window of previous pivotal tokens in generation and neglects the extensive input that contains essential information for generating accurate responses. This oversight could lead to an inability to extract detailed information from prompts.
In summary, existing compression methods merely address the challenges encountered in real-world applications, such as document processing and multi-round chats, where prompts are exceptionally long yet require accurate information retrieval. In common use cases, the generated outputs, like summaries, code pieces, or retrieved data, are significantly shorter compared to the extensive input sequences from novels, entire code bases, or annual financial reports. Although these techniques may effectively reduce the KV cache size during the generation phase, they do not tackle the primary overhead and challenges arising from a lack of comprehension of complex input contexts, thus leaving the critical issues unresolved.
@@ 3 Observations
In this section, we present our observations regarding the patterns in the Query-Key matrix during token generation. We discuss how these patterns can be potentially exploited for KV cache compression. Our findings are based on the analysis of various generation contexts and the behavior of attention mechanisms in LLMs and are concluded into three key observations as follows:
  1. 1.
    Pattern consistency across contexts: Irrespective of the generation context length, we observed that specific keys within the prompt consistently exhibit higher attention weights. Such “active” keys tend to follow stable patterns that appear to be intrinsically related to the structure and content of the prompt. (Sec. 3.1)
  2. 2.
    Invariance to question positions in summarization tasks: In the context of long summarization and question-answering tasks, the positioning of questions within the prompt (either at the beginning or the end) does not significantly alter the consistency of attention patterns observed. This suggests a level of robustness in how we can obtain the attention of relevant features trivially, regardless of the position of questions. (Sec. 3.2.1)
  3. 3.
    Contextual dependency of patterns: The observed attention patterns are highly context-sensitive, indicating a strong association with the specific instructions posed by the user (Sec. 3.2.2). Thus, a context-aware KV compression approach can potentially lead to better performance.
To structure our experimental analysis coherently, we introduce the following terminologies:
Prompt Length ($L_{\text{prompt}}$): The total length of the user-provided input. Prefix Length ($L_{\text{prefix}}$): The length of the input preceding the observation window. It is part of the prompt and does not include the observation window. Observation Window ($L_{\text{obs}}$): The last segment of the prompt. This window is crucial for analyzing the influence of different contexts on attention patterns. These definitions are interconnected as follows:
@@
Voting: The process of calculating attention weights for each query within the observation window across all heads, aggregating these weights to highlight the prefix positions that are considered most significant.
For a single batch of sequence, formally:
@@
$I$ $=\text{Top}_{k}(\mathbf{C},k)$ (3)
where $\text{Top}_{k}(\mathbf{T},k)$ selects the indices of the top $k$ values in tensor $\mathbf{T}$ per head, $k$ is defined as $\left\lfloor\textit{p}\times L_{\text{prefix}}\right\rfloor$. The tensor $\mathbf{W}_{\text{obs}}\in\mathbb{R}^{N\times L_{\text{obs}}\times L_{\text{prefix}}}$ represents the subset of the prompt softmax-normalized attention features over $N$ heads.
Hit Rate: The hit rate, $H$, quantifies the effectiveness of the voting mechanism by measuring the ratio of attention features identified as significant by the voting process that are also essential in the generation outcome, calculated as:
@@
$\mathbf{A}_{\text{cur}}\in\mathbb{R}^{N\times L_{\text{prefix}}}$ represents the attention features between the current generated query and prefix keys. The threshold operation filters $\mathbf{A}_{\text{cur}}$ to retain only values exceeding $\theta$, indicating significant attention activations. The overlap $\mathbf{O}$ between these significant activations and the mask $\mathbf{M}$ quantifies the alignment of the current attention with previously identified significant features. The hit rate $H$ is then computed as the ratio of the sum of overlap $\mathbf{O}$ to the sum of significant activations $\mathbf{A}_{\text{threshold}}$, providing a metric for the efficacy of the attention mechanism in recognizing and emphasizing important attention features within the context. We can use $\mathcal{H}(\mathbf{M}_{\text{threshold\_cur}},\mathbf{M}_{\text{vote\_obs}})$ denote combination of eq. 7 and eq. 8. We use $p=0.05$ (top 5% location per head) and $\theta=0.05$ (note it is a large value due to the softmax function over a long sequence) for the observation experiments. The model we probe is Mistral-7B-Instruct-v0.2.
3.1 Observations in Multi-Turn Conversations
Figure 2: The layer-wise average hit rate of important positions utilized along token generation with an average input length exceeding 3k.
Refer to captionRefer to caption
Figure 3: The layer-wise average hit rate of important positions utilized by prompts with questions at the beginning and the end.
This study examines if the positions of features identified as crucial in the observation window maintain their significance in the subsequent token generation. The analysis utilizes samples from Ultrachat [11], a multi-turns, high-quality instruction dataset consisting of 1.4 million dialogues. We further filter the sequences with response length greater than 512 and prompt length greater than 3k. In the experiment, we split the generated tokens into 4 context windows, each spanning 128 tokens, to compute the averaged hit rates of these windows versus the observation window with size 32. According to the findings presented in Fig.2, important keys in prefixes obtained from voting in observation windows exhibit remarkable consistency throughout the generation process, as evidenced by high hit rates.
3.2 Observations in Long Document QA
To further validate this finding, we also observe on multiple long documents QA datasets including QMSum [12], a query-based multi-domain meeting summarization; Openreview [13], a collection of papers from openreview.net; SPACE [14], an extractive opinion summarization in quantized transformer spaces.
3.2.1 Effectiveness of Instruction Positions
Our investigation also extends to the significance of instruction positioning on the interpretability of LLMs and their selection of important features. We calculate the average hit rate for the responses using the same observation window size of 32 as in the previous experiment. Our results shown in Fig. 3 indicate that across all three datasets, the hit rates are consistently high regardless of whether instructions are positioned before or after extensive supplementary contexts. This consistency suggests that the patterns identified by observation windows are independent of the question’s positions.
3.2.2 Effectiveness of Various Instructions for One Document
Furthermore, we investigate whether instructions will affect the selection of important features even if the provided context is the same. Our experiment utilizes different instructions on the same document and selects the important features based on the observation window that consists of both the instructions and their corresponding responses. Then we calculate the hit rates between important features selected by different instruction-response pairs within the same document by using $\mathcal{H}(\text{M}_{\text{vote\_A}},\text{M}_{\text{vote\_B}})$. By varying the instructions, we observe that different instructions prioritize different prefix keys, as indicated by the descending trend in hit rates shown in Fig. 4.
Our findings reveal an interesting aspect of KV cache management in LLMs: the important attention features change with different instructions. This variability challenges the effectiveness of static compression methods that depend on constant weighted importance or fixed policies [7, 6, 8]. Thus, the complex relationship between context and related KV cache emphasizes the need for context-aware compression strategies and highlights the limitations of current methods that ignore this dynamic.
4 SnapKV
4.1 Basic Method
In the attention mechanism, keys and values are tensors containing information from the previous context. The linear growth in prompts will lead to exponential time complexity for generation due to the Query-Key matrix multiplication. SnapKV addresses this by keeping prompt KV cache counts constant during generation, significantly reducing serving times for long-context LLMs.
@@ 4.1 Basic Method
The fundamental approach of SnapKV involves identifying and selecting the most crucial attention features per head to create the new KV cache. SnapKV operates through two stages as shown in Fig. 1:
@@
  • Voting for Important Previous Features By the voting process defined previously (Eq. 1), we select the important features based on the observation window—defined as the last segment of the prompt. Sec. 3.1 highlights the consistency of these attention features throughout the sequence, suggesting that these features are vital for subsequent generation. Besides, we implement clustering to retain the features surrounding the selected features (See Sec.4.2).
@@
  • Update and Store Truncated Key and Value We concatenate these selected features with the window features, which encompass all features containing prompt information. We store back the concatenated KV caches for later use in generation and save the memory usage.
@@
12 attn_weights_sum = attn_weights[..., -window_size:, :-window_size].sum(dim=-2)
@@
14 attn_cache = pool1d(attn_weights_sum, kernel_size=kernel_size, padding=kernel_size//2, stride=1)
15 # Select top-k indices per head based on the pooled weights to identify important positions.
16 indices = attn_cache.topk(max_capacity_prompt - window_size, dim=-1).indices

### v2

# SnapKV: LLM Knows What You are Looking for Before Generation

@@ Abstract
We discover that each attention head in the model consistently focuses on specific prompt attention features during generation. Meanwhile, this robust pattern can be obtained from an ‘observation’ window located at the end of the prompts. Drawing on this insight, SnapKV automatically compresses KV caches by selecting clustered important KV positions for each attention head. Our approach significantly reduces the growing computational overhead and memory footprint when processing long input sequences. Specifically, SnapKV achieves a consistent decoding speed with a 3.6x increase in generation speed and an 8.2x enhancement in memory efficiency compared to the baseline when processing inputs of 16K tokens. At the same time, it maintains comparable performance to the baseline models across 16 long sequence datasets. Moreover, SnapKV can process up to 380K context tokens on a single A100-80GB GPU using HuggingFace implementation with minor changes, exhibiting only a negligible accuracy drop in the Needle-in-a-Haystack test. Further comprehensive studies suggest SnapKV’s potential for practical applications.
@@ 1 Introduction
Many leading LLMs have started to handle longer contexts, overcoming the difficulties in context maintenance and attention mechanism scalability, such as GPT-4 [1] and Command-R [2] with context length 128K, Claude-3 [3] with 200K, and Gemini-Pro-1.5 with 1M [4]. Despite their impressive capabilities, LLMs still face significant challenges when dealing with long context prompts. Specifically, the KV cache in attention calculation becomes less efficient when processing long context. During inference time, as prompt length increases, the decoding latency per step grows linearly due to the attention calculation across past KVs. Moreover, the large KV cache requires significant memory capacity, increasing hardware demands and limiting model scalability.
There are many approaches to mitigate these problems, such as KV cache eviction during generation stage  [5, 6, 7, 8]. However, most of these methods lack a detailed evaluation in long-context settings. Moreover, they mainly focus on compressing the KV cache appended during decoding steps, while overlooking the realistic problem of compressing KV cache for prompts, which is typically the bottleneck in memory efficiency. In practical applications like chatbots and agents, where prompts range from multi-turn conversations to extensive articles or codebases [1, 9, 10], prompts are often much larger than generated responses such as summaries and code pieces, thus creating significant inference latency and memory utilization overhead. Additional challenge lies in compressing KV cache for such vast prompts without losing crucial information for accurate generation, especially in scenarios with various noisy contexts.
Figure 1: The graph shows the simplified workflow of SnapKV, where the orange area represents the cluster of features per head selected by SnapKV. These features are then used to form new Key-Value pairs concatenated with the features in the observation window. Together, the selected prefix and observation windows constitute the new KV cache utilized for the generation.
In our paper, we find an important attention allocation phenomenon: only a portion of prompt tokens convey essential information for response generation, and these tokens remain unchanged during generation. To validate the robustness of this finding, we design a thorough set of experiments across diverse prompts in terms of length, format, and content. From our observations, we derive an innovative and intuitive method, SnapKV, which can smartly identify the attention allocation pattern and compress the KV cache for long sequence prompts without compromising the model’s accuracy. With its comprehensive design, SnapKV demonstrates its effectiveness on various datasets and can be easily integrated into popular deep-learning frameworks with just a few code adjustments. Our contributions are as follows:
  • We design experiments to explore the attention allocation pattern during generation, focusing on two key questions:
@@
      Is there a consistent attention allocation pattern for input sequence tokens?
@@
      Is it feasible to identify this pattern prior to the generation stage?
    Our finding suggests that for LLMs, the attention allocation of most input sequence tokens stay consistent during generation. Thus, LLMs knows what you are looking for before generation.
  • Inspired by our observations above, we develop an efficient and fine-tuning-free algorithm, SnapKV, which efficiently identifies critical attention features and compresses KV cache correspondingly with minimal model modification (See Fig. 1).
  • We evaluate SnapKV across diverse LLMs and long-sequence datasets. SnapKV shows comparable accuracy with full KV caching method while achieving improved decoding speed and memory efficiency. Meanwhile, we conduct the pressure test with Needle-in-a-Haystack to further demonstrate its memory efficiency and information retrieval ability.
@@ 2 Related Works
Many previous works compress the KV cache by selectively dropping KVs using different algorithms. In StreamLLM [5], only the most recent tokens and attention sinks (first few tokens) are retained to reduce the KV cache size, making it lose the important information carried by the discarded middle tokens  [^1] https://github.com/mit-han-lab/streaming-llm?tab=readme-ov-file#faq. Heavy-Hitter Oracle (H2O) [6] introduces a policy that greedily drops KVs during generation based on a scoring function derived from cumulative attention. While this approach effectively compresses the KVs appended to the cache during generation, it overlooks compression of prompt KVs, which is crucial for reducing memory and computational overhead. Building on a similar concept, Adaptive KV Compression (FastGen) [8] implements a dual-phase algorithm that encompasses four KV cache compression policies. Initially, it identifies optimal policies through profiling results obtained from prompt encoding. Subsequently, it dynamically evicts caches during the generation phase based on these policies. Nonetheless, it faces the similar problem with H2O. ScissorHands [7] focuses on identifying and retaining pivotal tokens that exhibit a consistent attention weight pattern with previous token windows during generation steps. However, this method concentrates solely on the window of previous pivotal tokens in generation and neglects the extensive prompt that contains essential information for generating accurate responses. This oversight could lead to an inability to extract detailed information from prompts.
In summary, existing methods have not effectively addressed the challenges encountered in real-world applications, where prompts are exceptionally long yet require accurate information retrieval. Although these techniques may reduce the KV cache size during generation, they do not address the primary challenges of understanding complex prompt contexts, leaving critical issues unresolved.
Figure 2: The overlap rates between attention features of the input sequence, selected by various windows along the input and during generation, with each line representing a model layer.
Figure 3: The layer-wise overlap rates between input sequence attention features selected by the last window of input sequence and those selected by 4 windows along generation.
@@ 3 Observations
In this section, we present our observations regarding the attention allocation patterns in the Query-Key matrix during token generation. Our analysis utilizes samples from Ultrachat [11], a multi-turns, high-quality instruction dataset consisting of 1.4 million dialogues. We further filter the sequences with response length greater than 512 and prompt length greater than 3k. Our findings are concluded into two key observations as follows:
  • Pattern can be identified before generation. In this experiment, we split the attention features of input sequence of each layer into multiple windows, each with 128 tokens, and calculate the averaged attention weights of the last 20 windows separately. To understand the attention allocation patterns along input sequences, we calculate the overlap rates between important attention features of input sequence (those with high average attention weights) identified by each window and the actual ones used by generation. The experimental results are shown in Fig. 3. We observe that the last window of input sequence recognizes highly similar attention allocation pattern with the actual generation.
  • Pattern is consistent during generation. We study if the positions of features identified as crucial in the last window of input sequence maintain their significance in the subsequent token generation. In the experiment, we split the generated tokens into 4 windows for every layer, each spanning 128 tokens, to compute the averaged overlap rates of these windows versus the last window of input sequence. As shown in Fig. 3, active attention features of input sequence obtained from the last window exhibit remarkable consistency throughout the generation process, as evidenced by high overlap rates.
4 SnapKV
In the attention mechanism, the growth in prompts will significantly increase time complexity for generation due to the Query-Key matrix multiplication. SnapKV addresses this issue by maintaining a constant amount of prompt KVs during generation, significantly reducing serving times for long-context LLMs. To structure our method coherently, we propose the following terminologies:
  • Prompt Length ($L_{\text{prompt}}$): The total length of the user-provided input.
  • Observation Window ($L_{\text{obs}}$): The last segment of the prompt. This window is crucial for analyzing the influence of different contexts on attention allocation patterns.
  • Prefix Length ($L_{\text{prefix}}$): The length of the input preceding the observation window. It is part of the prompt and does not include the observation window. Overall, we have:
@@ 4 SnapKV
  • Voting: The process of calculating attention weights for each query within the observation window across all heads, aggregating these weights to highlight the prefix positions that are considered most significant. For a single batch of sequence, formally:
@@
    $\quad I=\text{Top}_{k}(\mathbf{C},k)$ (3)
    where $\text{Top}_{k}(\mathbf{C},k)$ selects the indices $I$ of the top $k$ values in tensor $\mathbf{C}$ per head. $k$ is defined as $\left\lfloor\textit{p}\times L_{\text{prefix}}\right\rfloor$, where $p$ stands for the compression rate. The tensor $\mathbf{W}_{\text{obs}}\in\mathbb{R}^{N\times L_{\text{obs}}\times L_{\text{prefix}}}$ represents the subset of the prompt softmax-normalized attention features over $N$ heads.
  • Hit Rate: We define attention features above a predefined threshold $\theta$ during generation as important features. The hit rate, $H$, is the number of important features successfully selected by the previous voting process over the total number of important features. $H$ quantifies the effectiveness of the voting mechanism and is calculated as follows:
@@
    $\mathbf{A}_{\text{cur}}\in\mathbb{R}^{N\times L_{\text{prefix}}}$ represents the attention features between the current generated query and prefix keys. $\mathbf{M}$ selects attention features by indices. The threshold operation filters $\mathbf{A}_{\text{cur}}$ to retain only features with values over $\theta$, indicating important attention activations. The $\mathbf{O}$ measures the overlap between attention features selected by $\mathbf{M}_{\text{threshold\_cur}}$ and $\mathbf{M}_{\text{vote\_obs}}$, quantifying the alignment of the current attention with previously identified important features. The hit rate $H$ is then computed as the ratio of the sum of overlap $\mathbf{O}$ to the sum of important features $\mathbf{M}_{\text{threshold\_cur}}$, providing a metric for the efficacy of the attention mechanism in recognizing and emphasizing important attention features within the context. We use $\mathcal{H}(\mathbf{M}_{\text{threshold\_cur}},\mathbf{M}_{\text{vote\_obs}})$ to denote combination of Eq. 7 and Eq. 8.
4.1 Observation Window-based Algorithm
@@
The core approach of SnapKV involves identifying and selecting the most crucial attention features per head to create the compressed KV cache. Listing LABEL:lst:pytorch shows the PyTorch-style pseudo code of SnapKV. Overall, SnapKV operates through two stages as follows:
@@
  • Vote for important previous features. By the voting process defined above (Eq. 2), we select the important attention features based on the observation window. Sec. 3 highlights the consistency of the attention allocation pattern within observation windows throughout the generation, suggesting that these selected attention features are also vital for subsequent generation. Furthermore, we implement clustering to retain the features surrounding the selected attention features (Sec. 4.3). Line 8-17 shows the pseudo code of the voting process.
@@
  • Update and store compressed keys and values. We concatenate the selected attention features with all features within the observation window, which encompasses all features containing the necessary prompt information. Line 18- 24 shows the compressing process. The concatenated KVs are stored for later use in generation, thereby saving memory usage.
@@
11 vote = attn_weights[..., -window_size:, :-window_size].sum(dim=-2)
@@
13 pool_vote = pool1d(vote, kernel_size=kernel_size, padding=kernel_size//2, stride=1)
14 # Select top-k indices based on the pooled weights to identify important positions.
15 indices = pool_vote.topk(max_capacity_prompt - window_size, dim=-1).indices
@@
4.2 Robustness Analysis of Hit Rate
To understand the robustness of the observation window-based algorithm, we analyze its hit rate on multiple long documents QA datasets including QMSum [12], a query-based multi-domain meeting summarization; Openreview [13], a collection of papers from openreview.net; SPACE [14], an extractive opinion summarization in quantized transformer spaces. The model we probe is Mistral-7B-Instruct-v0.2. Overall, we want to answer the following two questions:
  1. 1.
    Does the nature of instructions in the prompt affect the hit rate?
  2. 2.
    Does the context and instruction positioning affect the hit rate?
4.2.1 Contextual Dependency of Patterns
We analyze whether instructions will affect the selection of important features even if the provided context is the same. Our experiment utilizes different instructions on the same document and selects the important features based on the observation window that consists of both the instructions and their corresponding responses. Then we calculate the hit rates between important features selected by different instruction-response pairs within the same document by using $\mathcal{H}(\text{M}_{\text{vote\_A}},\text{M}_{\text{vote\_B}})$. By varying the instructions, we observe that different instructions prioritize different prefix attention features, as indicated by the descending trend in hit rates shown in Fig. 4. Our findings reveal an interesting aspect of KV cache in LLMs: the important attention features change with different instructions. This variability challenges the effectiveness of static compression methods that depend on constant weighted importance or fixed policies [7, 6, 8]. Thus, the complex relationship between context and related KV cache emphasizes the need for context-aware compression strategies and highlights the capability of SnapKV that recognizes this dynamic.
4.2.2 Invariance to Instruction Positions
Our analysis also extends to the significance of instruction positioning on the interpretability of LLMs and their selection of important features. We calculate the average hit rate for the responses using the same observation window size as in the previous experiment. Our results shown in Fig. 5 indicate that across all three datasets, the hit rates are consistently high regardless of whether instructions are positioned before or after extensive supplementary contexts. This consistency suggests that SnapKV is able to identify attention allocation patterns regardless of the question’s positions.
Figure 5: The layer-wise average hit rate of important positions used by prompts with questions at the beginning and the end.